OpenStudy (pulsified333):

The graph y=(7x−2)/e^x has a horizontal tangent at x=

jimthompson5910 (jim_thompson5910):

This is nitpicky, but it should say `has a horizontal tangent at y =` the `x=` should be `y=`

jimthompson5910 (jim_thompson5910):

anyways, look at a table of values. Let x get larger and larger (10 --> 100 ---> 1000, etc) and see what happens to y

jimthompson5910 (jim_thompson5910):

yes the denominator is growing much much faster than the numerator, so overall the fraction gets closer to 0

OpenStudy (anonymous):

Actually, I think the \(x=\) is fine, since they use the word "at" rather than "which is"

jimthompson5910 (jim_thompson5910):

not sure why you deleted that since you were right

OpenStudy (pulsified333):

delete what?

OpenStudy (pulsified333):

That's weird i never deleted a comment.

jimthompson5910 (jim_thompson5910):

oh my bad I misread the problem entirely I though it said asymptote

OpenStudy (pulsified333):

its okay lol

jimthompson5910 (jim_thompson5910):

what you'll need to do is take the derivative and then set it equal to zero, and solve for x

jimthompson5910 (jim_thompson5910):

you'll use the quotient rule

OpenStudy (anonymous):

We're looking for \(x_0\) such that:\[ \frac{dy}{dx}\Bigg|^{x=x_0} =0 \]

OpenStudy (pulsified333):

the derivative is e^-x(9-7x)

OpenStudy (anonymous):

Then \(9-7x=0\) would give a root.

OpenStudy (pulsified333):

so does that mean that x=9/7?

jimthompson5910 (jim_thompson5910):

h(x) = f(x)/g(x) let h(x) be set to (7x−2)/(e^x) so f(x) = 7x-2 g(x) = e^x

OpenStudy (pulsified333):

it worked :D thanks man

jimthompson5910 (jim_thompson5910):

f(x) = 7x-2 f ' (x) = 7 g(x) = e^x g ' (x) = e^x ---------------------------- h(x) = f(x)/g(x) h ' (x) = [f ' (x)*g(x) - f(x)*g ' (x)]/[g(x)]^2 h ' (x) = [7*e^x - (7x-2)*e^x]/[e^x]^2 h ' (x) = [e^x*(7 - 7x+2)]/[e^(2x)] h ' (x) = [e^x*(9 - 7x)]/[e^(2x)] h ' (x) = [9 - 7x]/[e^(x)] so you have the correct derivative

OpenStudy (pulsified333):

@jim_thompson5910 yeah :D